# Test-Time Scaling of Diffusion Models via Noise Trajectory Search

### v1

@@ Appendix C Search Algorithms
$(N,K)$-Zero-Order Search $NKT$
$(N,K)$-$\epsilon$-greedy $NKT$
Appendix D Hyperparameter Settings & Scaling
@@
Figure 3: Performance scaling with number of noise candidates. Performance of sampling methods as the number of noise candidates $N$ per timestep increases, measured on three reward functions (brightness, compressibility, ImageNet). $\epsilon$-greedy achieves the best scaling law, attaining the highest rewards at the majority of $N$ values; however, it experiences non-monotonic gains due to its greedy local selection, whereas MCTS—by approximating exhaustive search—shows steady, monotonic improvement.
D.1 Scaling with Increasing $N,K$
@@
D.2 Sweeps over $\lambda,\epsilon$
@@
Appendix E Qualitative Results
@@
  • and (CLIP) emphasizing distinctive features of the image subject. Examples (see Figure 7) include the curved beak and wide brown region around the eyes of the kākāpō and the unique red nose of the red wolf. It’s particularly remarkable that just noise optimization alone for the CLIP reward increases faithfulness to elements of the prompt like “three rockhoppers in a line;” following detailed prompts like these is a notorious and long-standing problem in diffusion [Black et al., 2024, Kim et al., 2025], and none of these features together were captured by the other sampling methods besides $\epsilon$-greedy.
@@
Appendix F Broader Impacts
@@
Appendix G Limitations and Future Work
@@ G.1 Compute Efficiency
To mitigate this, recalling Figure 2, as a proof-of-concept, we run zero-order search and $\epsilon$-greedy with $K=20$ for only $\{t:0.01\leq\sigma_{t}\leq 1\}$, $K=1$ otherwise. This yields rewards within ±$\pm 0.04$ of the original results (where $K=20$ is used at all denoising steps), but cuts the NFEs by more than half. We recognize exploration of similar approaches to increase computational efficiency as an important direction for future research.
G.2 Population-Based Metrics
@@ G.4 Future Work
  • Reducing the computational overhead of per-step search, perhaps via adaptive scheduling of NFEs or learned proposal distributions
  • Exploring alternative bandit solvers or hybrid tree-bandit strategies to improve sample efficiency
  • Incorporating regularization strategies to mitigate reward overoptimization
  • More sophisticated searches that take into account the interaction between different diffusion steps (e.g., formulating correlation/dependencies between noise candidates across timesteps)
  • Learning compute-optimal regions during denoising at which to search
  • Using ensembles of rewards, e.g. average of several VLM reward metrics
  • Extending to human-in-the-loop settings where only preferences/ratings and no closed-form reward is provided

### v2

@@ Appendix C Search Algorithms
$(N,K)$-Zero-Order Search NKTNKT
$(N,K)$-$\epsilon$-greedy NKTNKT
Appendix D $\epsilon$-greedy Regret Analysis
To strengthen our claims on $\epsilon$-greedy’s superiority, we derive its regret bound, showing that—even with only black-box access to a Lipschitz surrogate—simple regret decays sublinearly in the total number of global samples.
Theorem 1 (Regret of $\epsilon$-greedy search)
Let the noise at diffusion step $t$ be $\mathbf{z}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I}_{d})$. Fix a confidence level $\eta$ and truncate noise-space to the high-probability ball $\Gamma_{t}=\left\{\mathbf{z}\in\mathbb{R}^{d}:\lvert\mathbf{z}\rvert\leq\sigma_{t}\sqrt{d\ln(1/\eta)}\right\}.$ Assume the surrogate reward $\tilde{R}_{t}:\Gamma_{t}\to[0,1]$ is $L$-Lipschitz and satisfies
$\left|\tilde{R}_{t}-r(\mathbf{x}_{0}(\mathbf{z}),\mathbf{c})\right|\leq\delta\qquad\forall\>\mathbf{z}\in\Gamma_{t}.$
Run Algorithm 1 ($\epsilon$-greedy local search) for $K$ iterations with $N$ candidate points per iteration and global-exploration probability $\epsilon$, and let the final pivot be $\mathbf{p}_{t}$. Then the expected simple regret at step $t$ obeys
$\mathbb{E}\!\left[\max_{\mathbf{z}\in\Gamma_{t}}\widetilde{R}_{t}(\mathbf{z})-\widetilde{R}_{t}(\mathbf{p}_{t})\right]\leq\delta+C_{d}\bigl{(}\epsilon NK\bigr{)}^{-1/d},$
where the constant $C_{d}>0$ depends only on the dimension $d$. Summing over $T$ diffusion steps yields
$R_{\text{tot}}=\sum_{t=1}^{T}\mathbb{E}\!\left[\max_{\mathbf{z}\in\Gamma_{t}}\widetilde{R}_{t}(\mathbf{z})-\widetilde{R}_{t}(\mathbf{p}_{t})\right]\leq T\left[\delta+C_{d}(\epsilon NK)^{-1/d}\right].$
Proof.
By Gaussian concentration, $\mathbf{z}_{t}$ lies in $\Gamma_{t}$ with probability at least $1-\eta$, so we may restrict attention to $\Gamma_{t}$ at negligible cost. Each global-exploration draw is uniform on $\Gamma_{t}$, and in total the algorithm collects $M=\epsilon NK$ such draws. It is a well-known result that the best of $M$ uniform samples over a $d$-dimensional, Lipschitz-reward domain achieves expected simple regret on the order of $M^{-1/d}$. Since the $\epsilon$-greedy pivot never decreases in surrogate reward, its final surrogate value is at least that of the best global sample, inheriting the same $\Theta(M^{-1/d})$ rate. Finally, replacing the surrogate by the true reward adds at most $\delta$, yielding the per-step bound; summing over $T$ steps gives the stated total regret. ∎
Note the following key points:
  • Polynomial decay in global samples. With $M=\epsilon NK$ global draws per step, regret scales as $O(M^{-1/d})$, so increasing the total sample budget drives simple regret to zero at the classic random-search rate.
  • Bias floor $\delta$. In practice we can take $\delta$ to be quite low as it is simply the approximation error induced by application of Tweedie’s formula, $|r(\mathbf{x}_{0},\mathbf{c})-r(\hat{\mathbf{x}}^{(t)}_{0},\mathbf{c})|$.
Appendix E Visualizing $\epsilon$-greedy Noise Search
We display a animation visualizing the $\epsilon$-greedy search process at this link. For each timestep $t$, over the $K=20$ local search iterations, we plot the 2D t-SNE [van der Maaten and Hinton, 2008] projection of the pivot and track it over time. The animation displays visually the emergent properties of $\epsilon$-greedy we describe in subsubsection 5.1.3 of the main paper: at early and late denoising steps, the search is mostly random global exploration; but at intermediate steps, after a couple initial random draws, all subsequent iterations are local hill-climbing.
Appendix F Additional Experiments
F.1 Scaling up dataset size
Below, we report EDM results using the ImageNet reward on $300$ images, a $10\times$ increase in dataset size. This confirms the superiority of our $\epsilon$-greedy search.
Table 4: EDM results by sampling method using classifier reward. We use the same algorithm hyperparameters as in Table 1 of the main paper.
Reward Naive Best-of-4 Sampling Beam Search MCTS Zero-Order Search $\bm{\epsilon}$-greedy
Classifier $0.3738$ $0.6413$ $0.6928$ $0.7311$ $0.6818$ $\mathbf{0.9790}$
F.2 Increasing task complexity
We provide results evaluating methods on a far harder set of prompts containing compositional information Black et al. [2024], specifically those of the form “a [SUBJECT] [PHRASE],” e.g. “a lion washing the dishes.” Table 5 displays these results; $\epsilon$-greedy is again superior across the board.
Table 5: SD results by sampling method, across reward functions, on compositional prompts. We use the same hyperparameters as in Table 2 of the main paper.
Method Brightness Compressibility CLIP
Naive Sampling $0.4579$ $0.6866$ $0.2578$
Best-of-4 Sampling $0.5257$ $0.7639$ $0.2618$
Beam Search $0.5179$ $0.7690$ $0.2646$
MCTS $0.5921$ $0.7717$ $0.2954$
Zero-Order Search $0.4573$ $0.7128$ $0.2802$
$\bm{\epsilon}$-greedy $\mathbf{0.6312}$ $\mathbf{0.7869}$ $\mathbf{0.3053}$
F.3 Increasing reward complexity
To test our method on harder rewards capturing more fine-grained details of images, we provide results evaluating each method using the counting reward from [Kim et al., 2025b], normalized to $[0,1]$. Each prompt is a comma-separated list of 1166 clauses, each of the form “[NUMBER] [OBJECT],” e.g. “five horses, three cars, one train, five airplanes.” The counting reward function leverages open-source object detection models to compute the MSE between the actual and ground truth number of objects in the generated image, averaged across all objects described in the prompt. Even with this hard, non-differentiable counting reward function, we again see that $\epsilon$-greedy is superior across the board.
Table 6: SD results by sampling method, using counting reward.
Reward Naive Best-of-4 Beam Search MCTS Zero-Order Search $\bm{\epsilon}$-greedy
Counting Reward $0.3076$ $0.3156$ $0.4318$ $0.4377$ $0.3846$ $\mathbf{0.5384}$
F.4 Varying model family, reward function
To support our claim of $\epsilon$-greedy’s superiority, we display results using a larger text-to-image model (SDXL [Podell et al., 2023]) and the text-to-image human preference reward model ImageReward [Xu et al., 2023a]. We use the same prompts and algorithm hyperparameters as in Table 2 of the main paper, and normalize the ImageReward score to $[0,1]$. Results are shown in Table 7.
$\epsilon$-greedy remains the highest-performing sampling method in this experimental setting as well, highlighting its superiority across reward functions and diffusion model families/sizes.
Table 7: SDXL results by sampling method, using ImageReward [Xu et al., 2023a] as the reward function.
Reward Naive MCTS Zero-Order Search $\bm{\epsilon}$-greedy
Image-Reward $0.4668$ $0.5152$ $0.5036$ $\mathbf{0.5830}$
F.5 Tweedie estimates vs. learned reward models, few-step models
A possible concern with our proposed approach is that predicting $\mathbf{x}_{0}$ solely using Tweedie’s formula with a base model remains too crude, potentially yielding relatively low benefits, especially during the initial sampling stages. Alternatives include using a learned soft value function on latents or a few-step model distilled from the corresponding base model, which might lead to better results.
We test this empirically, showing that using a Tweedie estimate is extremely competitive with, and often actually outperforms both these alternatives, across model architectures, sampling methods, and reward functions.
First, in the class-conditional image generation setting, we compare running the ImageNet classifier on the Tweedie estimate $\hat{x}_{0}^{(t)}$, versus running a classifier on the noisy latent $x_{t}$ itself at each timestep $t$. For this “noisy classifier” we adopt the ImageNet-1k $64\times 64$ classifier from Dhariwal and Nichol [2021], trained to classify latents across timesteps and noise scales. As with the regular classifier reward, the reward here is the noisy classifier’s probability assigned to the ground-truth class. Table 8 displays these results.
Table 8: EDM results by sampling method using either a classifier reward on the Tweedie estimate, or a learned classifier on the noisy latents.
Reward Beam Search Zero-Order Search $\bm{\epsilon}$-greedy
ImageNet classifier on Tweedie-estimate $0.9666$ $\mathbf{0.6261}$ $0.9885$
Learned ImageNet classifier on latents $\mathbf{0.9792}$ $0.6099$ $\mathbf{0.9998}$
Second, in the harder text-to-image generation setting, we compare computing the CLIP reward on the Tweedie estimate, with instead computing the CLIP reward on the clean sample generated using a one-step distilled model from the current latent and noise candidate, at each timestep $t$. As in the main paper, we use Stable Diffusion v1.5, and for the one-step model use InstaFlow-0.9B [Liu et al., 2024] which was distilled from SD1.5. Results are shown in Table 9.
Table 9: SD results by sampling method using CLIP reward either on the Tweedie estimate or the predicted clean sample using a one-step distilled model.
Reward Beam Search Zero-Order Search $\bm{\epsilon}$-greedy
Tweedie-estimate $\mathbf{0.2682}$ $0.2689$ $\mathbf{0.2973}$
One-step model distilled from SD1.5 $0.2671$ $\mathbf{0.2825}$ $0.2820$
Appendix G Hyperparameter Settings & Scaling
@@
Figure 3: Performance scaling with number of noise candidates. Performance of sampling methods as the number of noise candidates $N$ per timestep increases, measured on three reward functions (brightness, compressibility, ImageNet). $\epsilon$-greedy achieves the best scaling law, attaining the highest rewards at the majority of NN values; however, it experiences non-monotonic gains due to its greedy local selection, whereas MCTS—by approximating exhaustive search—shows steady, monotonic improvement.
G.1 Scaling with Increasing $N,K$
@@
G.2 Varying $K$ across timesteps
Recall that $\epsilon$-greedy requires $NKT$ NFEs for $K$ local search iterations. Note, however, that this assumes $K$ is fixed across timesteps. Letting $K_{t}$ be the number of local search iterations at timestep $t$, we can significantly bring down $\mathbb{E}_{t}[K_{t}]$ by noting that $K_{t}$ only need be high at intermediate denoising steps where de-mixing occurs, and can be low at extreme (beginning and end) timesteps. We report $\epsilon$-greedy performance as a function of $\mathbb{E}_{t}[K_{t}]$ below in Figure 4.
Figure 4: EDM results by sampling method, varying $\mathbb{E}_{t}[K_{t}]$. We use the same class labels and (other than $K$) algorithm hyperparameters as in Table 1 of the main paper.
Indeed, $\epsilon$-greedy attains strong performance across all reward functions, even with extremely low average $K_{t}$ (can retain performance within $0.002$ of the highest value with $\mathbb{E}_{t}[K_{t}]$ as low as 4), highlighting our method’s computational efficiency compared to MCTS.
G.3 Sweeps over $\lambda,\epsilon$
@@
G.4 Varying classifier-free guidance scale, number of timesteps $T$
See Table 10 and Table 11 above. These results confirm the superiority of $\epsilon$-greedy irrespective of guidance scale and $T$.
Table 10: SD results by sampling method, varying the classifier-free guidance scale. We evaluate only CLIP reward as the brightness and compressibility rewards are generally insensitive to classifier-free guidance.
Method $\mathbf{cfg=1.5}$ $\mathbf{cfg=3.0}$ $\mathbf{cfg=7.5}$ $\mathbf{cfg=9.0}$ $\mathbf{cfg=12.0}$
Naive Sampling $0.2446$ $0.2635$ $0.2457$ $0.2650$ $0.2671$
Best-of-4 Sampling $0.2612$ $0.2743$ $0.2776$ $0.2783$ $0.2759$
Beam Search $0.2673$ $0.2314$ $0.2682$ $0.2801$ $0.2780$
MCTS $0.2828$ $0.2712$ $0.2885$ $0.2971$ $0.2827$
Zero-Order Search $0.2539$ $0.2612$ $0.2689$ $0.2443$ $0.2801$
$\bm{\epsilon}$-greedy $\mathbf{0.2958}$ $\mathbf{0.3092}$ $\mathbf{0.2973}$ $\mathbf{0.3055}$ $\mathbf{0.2994}$
Table 11: EDM results by sampling method, varying the value of $T$. Each cell contains two values, corresponding to scores under the Brightness / Compressibility reward functions.
Method T=18{T=18} T=32{T=32} T=64{T=64} T=128{T=128} T=256{T=256}
Naive Sampling $0.4965/0.3563$ $0.3912/0.3921$ $0.4621/0.4549$ $0.4849/0.4052$ $0.4541/0.4250$
Best-of-4 Sampling $0.5767/0.4220$ $0.5666/0.4992$ $0.6475/0.5084$ $0.5801/0.5101$ $0.5923/0.4917$
Beam Search $0.6334/0.4679$ $0.6013/0.4321$ $0.5742/0.5023$ $0.6162/0.4517$ $0.6618/0.4879$
MCTS $0.7575/0.5395$ $0.6910/0.5153$ $0.6831/0.5624$ $0.6957/0.5018$ $0.7494/0.5789$
Zero-Order Search $0.6083/0.3751$ $0.9999/0.7417$ $99/\mathbf{0.7607} 0.9999/0$.76710.9999/0.7671 $0.9999/\mathbf{0.7677}$
$\bm{\epsilon}$-greedy $\mathbf{0.9813}/\mathbf{0.7208}$ $\mathbf{0.9999}/\mathbf{0.7424}$ $\mathbf{0.9999}/0.7588$ $\mathbf{0.9999}/\mathbf{0.7680}$ $\mathbf{0.9999}/0.7672$
Appendix H Qualitative Results
@@
  • and (CLIP) emphasizing distinctive features of the image subject. Examples (see Figure 8) include the curved beak and wide brown region around the eyes of the kākāpō and the unique red nose of the red wolf. It’s particularly remarkable that just noise optimization alone for the CLIP reward increases faithfulness to elements of the prompt like “three rockhoppers in a line;” following detailed prompts like these is a notorious and long-standing problem in diffusion [Black et al., 2024, Kim et al., 2025a], and none of these features together were captured by the other sampling methods besides $\epsilon$-greedy.
@@
Appendix I Broader Impacts
@@
Appendix J Discussion and Next Steps
@@
To mitigate this, recalling Figure 2, as a proof-of-concept, we run zero-order search and $\epsilon$-greedy with $K=20$ for only $\{t:0.01\leq\sigma_{t}\leq 1\}$, $K=1$ otherwise. This yields rewards within $\pm 0.04$ of the original results (where $K=20$ is used at all denoising steps), but cuts the NFEs by more than half. We recognize exploration of similar approaches to increase computational efficiency as an important direction for future research.
J.2 Population-Based Metrics
@@
Reducing computational overhead.
Future work should aim to lower the cost of per-step search, for instance through adaptive scheduling of NFEs or by learning proposal distributions that make candidate selection more efficient.
Exploring alternative bandit solvers.
Hybrid methods, such as combining tree-based search with bandit algorithms, may improve sample efficiency and accelerate convergence.
Mitigating reward overoptimization.
Regularization strategies could help reduce overfitting to specific reward signals.
Modeling diffusion-step interactions.
More sophisticated search approaches should account for correlations between noise candidates across timesteps, rather than treating each step independently.
Learning compute-optimal regions.
Identifying points during denoising where search is most valuable could allow resources to be allocated selectively, improving efficiency without sacrificing quality.
Using reward ensembles.
Averaging across multiple reward metrics, such as different VLM-based signals, can provide more stable guidance than relying on a single objective.
Human-in-the-loop extensions.
Future systems could be adapted to operate in preference-based settings where only human ratings or feedback are available instead of closed-form reward functions.